OpenStudy (anonymous):

please... find the line that satisfies the condition * y-intercept 6 , parallel to the line 2x+3y+4=0

OpenStudy (radar):

First solve for slope fromthe line equation that the problem has provided you 2x+3y+4=0

OpenStudy (radar):

The slope will be the coefficient of x when you have "rearranged" the formula with y isolated on the left and everything else to the right of the equal sign.

OpenStudy (radar):

Use rules of algebra for the "rearranging"

OpenStudy (anonymous):

thanks a lot

OpenStudy (radar):

Let me know if you have any more questions regarding this problem. I am here to help you learn how to solve.......not to provide answers for your homework.

OpenStudy (radar):

If anyone else is looking at this problem and is interested in its solution please post here.

OpenStudy (radar):

1. Determine slope: 2x+3y +4=0 Subtract 2x and 4 from both sides of equation 3y +2x - 2x +4 - 4 = -2x -4 "clean up" 3y=-2x-4 Divide both sides by 3 3y/3= -(2/3)x -4/3 "clean up" y =-(2/3)x -4/3 Coefficient of is -2/3 Slope = -2/3 2. Determine the line (rest of the equation) Equation will appear as: y = -(2/3) x + 6 Intercept is given in problem as 6 Parallel lines will have equal slope.

OpenStudy (anonymous):

thank you so much